You run a trail at 7 mph and walk back at 3 mph. If your total time is 1.5 hours, your round-trip distance is how many miles?
Answer by Alan3354(69443)

You run a trail at 7 mph and walk back at 3 mph. If your total time is 1.5 hours, your round-trip distance is how many miles?
-----------
Avg speed for the round trip = 2*7*3/(7+3) = 4.2 mi/hr
----
d = r*t = 4.2*1.5 = 6.3 miles RT